1992 the life expectancy of males in a certain county was 69.4 years, in 1998 it was 72.9 years, let E repersent the life expectancy in year, and t repersent the number of years since 1992.
E(T)=___t+____
E(11)=_______

Answer by RAY100(1637)

E(t) = 69.4 + ,5833t
.
where 69.4 = E(0),,,.5833 = (72.9-69.4)/6
.
check,,,,E(6) = 69.4 +6(.58833) = 72.9,,,,,ok
.
E(11) = 69.4 + 11(.58833) = 75.8
.